# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v3

@@ I.2 The comprehensive two-orbital study
The operators $c_{i\mu\alpha\sigma}$, $n_{i\mu\alpha}$, $\mathbf{S}_{i\mu\alpha}$ take the same meanings as those in model (1) except for an extra index $\alpha=x/z$ labeling the orbital, and $\hat{P}$ is a projection operator projecting out the double occupancy in the same orbital of all sites. Note that $\mathbf{S}_{i\mu\alpha}$ for each orbital is spin-$\frac{1}{2}$ operator. $\epsilon_{\alpha}$ denotes the on-site energy of the orbital $\alpha$. We adpot the tight-binding (TB) parameters reported in Ref. [Daoxin_Yao2025], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The superexchange interactions are obtained through $J_{\parallel}\approx 4t_{\parallel}^{2}/U$ and $J_{\perp}\approx 4t_{\perp}^{2}/U$, with $U=10t_{\parallel}$. Finally $\varepsilon$ denotes the voltage between the two layers. Here, due to the weak super-exchange interaction between the $d_{z^{2}}$ orbitals in the layer, we do not consider this term in our model. In addition, the Hund’s coupling $J_{H}$ of La3Ni2O7 is generally considered to be in the range of 0.7 eV to 1 eV in past studies [lechermann2023, ouyang2023hund, WuWei2023charge], which only slightly larger than the largest hopping parameter $t_{\perp}=0.503~\mathrm{eV}$ and thus does not satisfy the premise of the Schriffer-Wolf transformation or perturbation theory, we do not apply it here. More details are provided in METHODS.
@@
Our SBMF results of Eq. (2) (see METHODS and the SI [SM]) are shown in Fig. 5. Fig. 5(a) shows the $\varepsilon$-dependence of the hole densities $\delta_{\mu\alpha}$. Obviously, the $\delta_{tz}$ enhances obviously with $\varepsilon$, suggesting that the top-$3d_{z^{2}}$ orbital is donating electrons. These donated electrons flow to the $3d_{x^{2}-y^{2}}$ orbitals in both layers, with more of them flowing to the bottom layer when $\varepsilon>0.1\text{ eV}$ while about half of them flow to the bottom layer when $\varepsilon\leq 0.1\text{ eV}$. Fig. 5(b) shows the $\varepsilon$-dependence of the pairing symmetry and the pairing gap amplitude of the bottom-layer $3d_{x^{2}-y^{2}}$ orbital. At low $\varepsilon\leq 0.03$ eV, the pairing symmetry is $s$-wave, whose pairing configuration is shown in Fig. 5(c), wherein the $3d_{z^{2}}$-orbital form interlayer $s$-wave pseudo-gap, while the $3d_{x^{2}-y^{2}}$ orbital form $s$-wave SC with coexisting intralayer and interlayer pairing. In this regime the interlayer pairing is suppressed by the enhancement of $\varepsilon$ while the intralayer pairing is enhanced. When $\varepsilon=0$, the interlayer pairing gap is the largest. When $\varepsilon$ is about $0.01\sim 0.03\text{ eV}$, the intralayer pairing gap is slightly larger than the interlayer pairing gap. When $\varepsilon>0.03$ eV, the pairing symmetry is dd($3d_{x^{2}-y^{2}}$)+iss($3d_{z^{2}}$), whose pairing configuration is shown in Fig. 5(d). In this state, the $3d_{z^{2}}$ orbital form interlayer $s$-wave pseudo-gap, while the bottom-layer $3d_{x^{2}-y^{2}}$ orbital form intralayer $d$-wave SC. When $\varepsilon$ enhances in this regime, the pairing amplitude for the $d$-wave part enhances promptly. For $\varepsilon>0.13$ eV, the pairing amplitude can go beyond $0.02$ eV. Then from the relation $T_{c}\approx 0.42\tilde{\Delta}$ for the $d$-wave SC illustrated in Fig. 3(b), we have got HTSC with $T_{c}\gtrsim 80$ K!
@@
II Discussion
In conclusion, we propose that an imposed strong perpendicular electric field can drive HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single-bilayer film of La3Ni2O7 at AP. The reason lies in that under the strong electric field, the electrons in the layer with higher potential energy will flow to the layer with lower potential energy, to fill the $3d_{x^{2}-y^{2}}$ orbitals in the latter layer. With considerably enhanced filling fraction, the $3d_{x^{2}-y^{2}}$ electrons in that layer just mimic the cuprates, which form intralayer $d$-wave HTSC with strongly enhanced $T_{c}$. Our combined one-orbital and two-orbital studies consistently verify this idea.
@@ II Discussion
III Methods
III.1 The one-orbital model
The SBMF theory is used to solve the one-orbital model (1). In the SBMF approach, the superexchange terms are decomposed in $\chi-\Delta$ channel, e.g. $\mathbf{S}_{it}\cdot\mathbf{S}_{ib}=-\frac{3}{8}\left(\left\langle\chi^{\perp{\dagger}}\right\rangle\chi^{\perp}_{i}+\mathrm{h.c.}+\left\langle\Delta^{\perp{\dagger}}\right\rangle\Delta^{\perp}_{i}+\mathrm{h.c.}\right)$, $\chi$ and $\Delta$ represents hopping and pairing operators respectively. These MF parameters are further solved in a self-consistent manner. The specific steps can be referenced from prior work [lu2023bilayertJ, kotliar1988, Lu2024interplay] and SI [SM].
We also employ the DMRG method [white1993dmrg, weng1999dmrg] to solve the ground state of the Hamiltonian (1) as a comparison for the SBMF approach. The tensor libraries TensorKit [jutho_2024_13950435] and FiniteMPS [Li_FiniteMPS_jl_2024] provide an implementation of the required symmetry [WEICHSELBAUM20122972, PhysRevResearch.2.023385]. We study the model on a $2\times 2\times L_{x}$ lattice with the open boundary conditions in the $x$ direction and choose $L_{x}=64$ for calculations. The matrix product state is constructed as shown in Fig. 6. We keep up to $D=12000$ $\mathrm{U}(1)_{\text{charge}}\times\mathrm{SU}(2)_{\text{spin}}$ multiplets in DMRG simulations and ensure the convergence accuracy of $10^{-6}$.
@@ III.1 The one-orbital model
The dominant pairing channel is related to pairing symmetry. For the case where interlayer pairing dominates, the pairing symmetry is restricted to $s$-wave pairing; while for the case where intralayer pairing in the bottom-layer dominates, we determine the pairing symmetry by the sign function $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]$. If $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]=-1$ holds for all $r$, the ground state can be identified as the $d$-wave pairing SC state. See SI [SM] for more details on DMRG.
III.2 The two-orbital model
@@
where $\delta_{\mu\alpha}=\left\langle{}b_{i\mu\alpha}b^{{\dagger}}_{j\mu\alpha}\right\rangle$ since holon condense at zero temperature. Under the electric field, we have $\delta_{bz}=0$ and $\delta_{tz}$, $\delta_{tx}$ and $\delta_{bx}$ are solved in a self-consistent manner by adjustment to onsite energies $\epsilon_{\alpha}$ (See SI [SM] for more details). The mean-field order parameters are represented by
@@
Notably, the spin-exchange $\tilde{J}_{\perp}$ of the Hamiltonian in Eq. (2) doesn’t produce a hopping term $\chi_{x}^{\perp}$ in Eq. (S22), which is the feature of such a bilayer system. Without interlayer hopping, a small interlayer spin-exchange $J_{\perp}$ leads to $\langle\chi^{\perp}\rangle\approx 0$.
@@
Appendix A A. Slave-boson mean-field treatment of the one-orbital model
@@
To get the knowledge of $\eta_{1}/\delta$ in different perpendicular electric field, we solve the tight-binding (TB) Hamiltonian with TB parameters reported in Ref. [Daoxin_Yao2025], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$, and an additional external electric field term $(\varepsilon/2)\sum_{i,\alpha,\sigma}(n_{it\alpha\sigma}-n_{ib\alpha\sigma})$. We plot the $(\delta,\eta_{1})$ points (Fig. A1), finding that $\eta_{1}/\delta\approx 1/2$. Thus we assume $\eta_{1}/\delta=1/2$ (i.e. $(n_{tx}-0.5)/(n_{bx}-0.5)=-\eta_{1}/\eta_{2}=-1/3$) in the one-orbital model study.
@@
We solve the Hamiltonian in Eq. (S1) by slave-boson mean-field (SBMF) theory [kotliar1988]. The electron operator ($c$) is decomposed into the product of a fermonic spinon operator ($f$) and a bosonic holon operator ($b$), i.e. $c^{{\dagger}}_{i\mu\sigma}=f^{{\dagger}}_{i\mu\sigma}b_{i\mu}$. Then the electron pairing operator is expressed as
@@
Appendix B B. Slave-boson mean-field treatment of the two-orbital model
@@
We adopt the TB parameters reported in Ref. [Daoxin_Yao2025], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The spin couplings are estimated as $J_{\parallel(\perp)}=\frac{4t_{\parallel(\perp)}^{2}}{U}$ with $U=10t_{\parallel}$ and $\tilde{J}_{\perp}=0.5J_{\perp}$. To limit the particle numbers of the two $E_{g}$ orbitals, we consider the Hamiltonian with Lagrange multipliers
@@
Appendix C C. More numerical simulation results by density matrix renormalization group
We apply the density matrix renormalization group (DMRG) approach [white1993dmrg, weng1999dmrg] to solve the ground state of Hamiltonian (S1) with $t_{\parallel}=1$, $J_{\parallel}=0.4t_{\parallel}$ and $\tilde{J}_{\perp}=0.8J_{\parallel}$. The definitions of all the operators and main results have been shown in the main text. Here we give more results.
@@
Appendix D D. The effect of interlayer coulomb interaction
@@
Appendix E Data availability
Appendix F Code availability
Appendix G Acknowledgement
Appendix H Author contributions
Appendix I Additional Information
Competing Interests
The authors declare no competing interests.

### v4

@@ II.2 The comprehensive two-orbital study
The operators $c_{i\mu\alpha\sigma}$, $n_{i\mu\alpha}$, $\mathbf{S}_{i\mu\alpha}$ take the same meanings as those in model (1) except for an extra index $\alpha=x/z$ labeling the orbital, and $\hat{P}$ is a projection operator projecting out the double occupancy in the same orbital of all sites. Note that $\mathbf{S}_{i\mu\alpha}$ for each orbital is spin-$\frac{1}{2}$ operator. $\epsilon_{\alpha}$ denotes the on-site energy of the orbital $\alpha$. We adpot the tight-binding (TB) parameters reported in Ref. [31], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The superexchange interactions are obtained through $J_{\parallel}\approx 4t_{\parallel}^{2}/U$ and $J_{\perp}\approx 4t_{\perp}^{2}/U$, with $U=10t_{\parallel}$. Finally $\varepsilon$ denotes the voltage between the two layers. Here, due to the weak super-exchange interaction between the $d_{z^{2}}$ orbitals in the layer, we do not consider this term in our model. In addition, the Hund’s coupling $J_{H}$ of La3Ni2O7 is generally considered to be in the range of 0.7 eV to 1 eV in past studies [55, 91, 126], which only slightly larger than the largest hopping parameter $t_{\perp}=0.503~\mathrm{eV}$ and thus does not satisfy the premise of the Schriffer-Wolf transformation or perturbation theory, we do not apply it here. More details are provided in METHODS.
@@
Our SBMF results of Eq. (2) (see METHODS and the SI [2]) are shown in Fig. 5. Fig. 5(a) shows the $\varepsilon$-dependence of the hole densities $\delta_{\mu\alpha}$. Obviously, the $\delta_{tz}$ enhances obviously with $\varepsilon$, suggesting that the top-$3d_{z^{2}}$ orbital is donating electrons. These donated electrons flow to the $3d_{x^{2}-y^{2}}$ orbitals in both layers, with more of them flowing to the bottom layer when $\varepsilon>0.1\text{ eV}$ while about half of them flow to the bottom layer when $\varepsilon\leq 0.1\text{ eV}$. Fig. 5(b) shows the $\varepsilon$-dependence of the pairing symmetry and the pairing gap amplitude of the bottom-layer $3d_{x^{2}-y^{2}}$ orbital. At low $\varepsilon\leq 0.03$ eV, the pairing symmetry is $s$-wave, whose pairing configuration is shown in Fig. 5(c), wherein the $3d_{z^{2}}$-orbital form interlayer $s$-wave pseudo-gap, while the $3d_{x^{2}-y^{2}}$ orbital form $s$-wave SC with coexisting intralayer and interlayer pairing. In this regime the interlayer pairing is suppressed by the enhancement of $\varepsilon$ while the intralayer pairing is enhanced. When $\varepsilon=0$, the interlayer pairing gap is the largest. When $\varepsilon$ is about $0.01\sim 0.03\text{ eV}$, the intralayer pairing gap is slightly larger than the interlayer pairing gap. When $\varepsilon>0.03$ eV, the pairing symmetry is dd($3d_{x^{2}-y^{2}}$)+iss($3d_{z^{2}}$), whose pairing configuration is shown in Fig. 5(d). In this state, the $3d_{z^{2}}$ orbital form interlayer $s$-wave pseudo-gap, while the bottom-layer $3d_{x^{2}-y^{2}}$ orbital form intralayer $d$-wave SC. When $\varepsilon$ enhances in this regime, the pairing amplitude for the $d$-wave part enhances promptly. For $\varepsilon>0.13$ eV, the pairing amplitude can go beyond $0.02$ eV. Then from the relation $T_{c}\approx 0.42\tilde{\Delta}$ for the $d$-wave SC illustrated in Fig. 3(b), we have got HTSC with $T_{c}\gtrsim 80$ K!
@@
III Discussion
In conclusion, we propose that an imposed strong perpendicular electric field can drive HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single-bilayer film of La3Ni2O7 at AP. The reason lies in that under the strong electric field, the electrons in the layer with higher potential energy will flow to the layer with lower potential energy, to fill the $3d_{x^{2}-y^{2}}$ orbitals in the latter layer. When the imposed electric field is weak, it acts as the “pseudo-Zeeman field” operating on the layer index which supresses the interlayer SC, possibly inducing the PDW state. With considerably enhanced filling fraction, the $3d_{x^{2}-y^{2}}$ electrons in that layer just mimic the cuprates, which form intralayer $d$-wave HTSC with strongly enhanced $T_{c}$. Our combined one-orbital and two-orbital studies consistently verify this idea.
@@ III Discussion
IV Methods
IV.1 The one-orbital model
The SBMF theory is used to solve the one-orbital model (1). In the SBMF approach, the superexchange terms are decomposed in $\chi-\Delta$ channel, e.g. $\mathbf{S}_{it}\cdot\mathbf{S}_{ib}=-\frac{3}{8}\left(\left\langle\chi^{\perp{\dagger}}\right\rangle\chi^{\perp}_{i}+\mathrm{h.c.}+\left\langle\Delta^{\perp{\dagger}}\right\rangle\Delta^{\perp}_{i}+\mathrm{h.c.}\right)$, $\chi$ and $\Delta$ represents hopping and pairing operators respectively. These MF parameters are further solved in a self-consistent manner. The specific steps can be referenced from prior work [80, 48, 81] and SI [2].
We also employ the DMRG method [125, 58] to solve the ground state of the Hamiltonian (1) as a comparison for the SBMF approach. The tensor libraries TensorKit [41] and FiniteMPS [68] provide an implementation of the required symmetry [123, 124]. We study the model on a $2\times 2\times L_{x}$ lattice with the open boundary conditions in the $x$ direction and choose $L_{x}=64$ for calculations. The matrix product state is constructed as shown in Fig. 6. We keep up to $D=12000$ $\mathrm{U}(1)_{\text{charge}}\times\mathrm{SU}(2)_{\text{spin}}$ multiplets in DMRG simulations and ensure the convergence accuracy of $10^{-6}$.
@@ IV.1 The one-orbital model
The dominant pairing channel is related to pairing symmetry. For the case where interlayer pairing dominates, the pairing symmetry is restricted to $s$-wave pairing; while for the case where intralayer pairing in the bottom-layer dominates, we determine the pairing symmetry by the sign function $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]$. If $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]=-1$ holds for all $r$, the ground state can be identified as the $d$-wave pairing SC state. See SI [2] for more details on DMRG.
IV.2 The two-orbital model
@@
where $\delta_{\mu\alpha}=\left\langle{}b_{i\mu\alpha}b^{{\dagger}}_{j\mu\alpha}\right\rangle$ since holon condense at zero temperature. Under the electric field, we have $\delta_{bz}=0$ and $\delta_{tz}$, $\delta_{tx}$ and $\delta_{bx}$ are solved in a self-consistent manner by adjustment to onsite energies $\epsilon_{\alpha}$ (See SI [2] for more details). The mean-field order parameters are represented by
@@
Notably, the spin-exchange $\tilde{J}_{\perp}$ of the Hamiltonian in Eq. (2) doesn’t produce a hopping term $\chi_{x}^{\perp}$ in Eq. (5), which is the feature of such a bilayer system. Without interlayer hopping, a small interlayer spin-exchange $J_{\perp}$ leads to $\langle\chi^{\perp}\rangle\approx 0$.
@@
V Data availability
VI Code availability
VII Acknowledgement
VIII Author contributions
IX Additional Information
Competing Interests
The authors declare no competing interests.
@@ IX Additional Information
X A. Slave-boson mean-field treatment of the one-orbital model
@@
To get the knowledge of $\eta_{1}/\delta$ in different perpendicular electric field, we solve the tight-binding (TB) Hamiltonian with TB parameters reported in Ref. [31], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$, and an additional external electric field term $(\varepsilon/2)\sum_{i,\alpha,\sigma}(n_{it\alpha\sigma}-n_{ib\alpha\sigma})$. We plot the $(\delta,\eta_{1})$ points (Fig. A1), finding that $\eta_{1}/\delta\approx 1/2$. Thus we assume $\eta_{1}/\delta=1/2$ (i.e. $(n_{tx}-0.5)/(n_{bx}-0.5)=-\eta_{1}/\eta_{2}=-1/3$) in the one-orbital model study.
@@
We solve the Hamiltonian in Eq. (S1) by slave-boson mean-field (SBMF) theory [48]. The electron operator ($c$) is decomposed into the product of a fermonic spinon operator ($f$) and a bosonic holon operator ($b$), i.e. $c^{{\dagger}}_{i\mu\sigma}=f^{{\dagger}}_{i\mu\sigma}b_{i\mu}$. Then the electron pairing operator is expressed as
@@
XI B. Slave-boson mean-field treatment of the two-orbital model
@@
We adopt the TB parameters reported in Ref. [31], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The spin couplings are estimated as $J_{\parallel(\perp)}=\frac{4t_{\parallel(\perp)}^{2}}{U}$ with $U=10t_{\parallel}$ and $\tilde{J}_{\perp}=0.5J_{\perp}$. To limit the particle numbers of the two $E_{g}$ orbitals, we consider the Hamiltonian with Lagrange multipliers
@@
XII C. More numerical simulation results by density matrix renormalization group
We apply the density matrix renormalization group (DMRG) approach [125, 58] to solve the ground state of Hamiltonian (S1) with $t_{\parallel}=1$, $J_{\parallel}=0.4t_{\parallel}$ and $\tilde{J}_{\perp}=0.8J_{\parallel}$. The definitions of all the operators and main results have been shown in the main text. Here we give more results.
@@
XIII D. The effect of interlayer coulomb interaction